The sides of a right triangle are a ,a + d,a +2d with a and d both positive .what is the value of ratio a :d.

Grade:10

1 Answers

Arun
6 years ago
 

Dear student

 

Solution: Applying Pythagoras Theorem,

 

we get, (a+2d)2=a2+(a+d)2

 

=> a2+4d2+4ad=a2+a2+d2+2ad

 

=> 3d2+2ad-a2=0

 

=> (d+a)(3d-a)=0

 

neglecting (d+a)=0, by solving for (3d-a)=0, we get a:d=3:1

 

Regards

Arun (askIITians forum expert)